A new 15-year mortgage had an initial balance of $123.8 thousand and an interest rate of 5.9%. What is the balance of this mortgage after 57 months? Assume that the borrower has made only the required payments. Express your answer in $ thousands, round to the nearest $0.01 thousand. E.g., if your answer is $125,573.78, record it as 125.57.
Answer #1

Initial Balance = $123.8 thousand

Interest Rate = 5.9%

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 123.8, N = 180, I = 0.059/12, FV = 0]

PMT = $1.04 thousand

Calculating Value of Loan after 57 months,

Using TVM Calculation,

FV = [PV = 123.8, PMT = -1.04, N = 57, I = 0.059/12]

FV = $95.50 thousand

Balance of mortgage after 57 months = $95.50 thousand